The volume of a circular disk varies as its thickness and the square of the diameter of its face. Two metallic disks having the thicknesses ¼ inch and 1/6 inch and the diameters of 5 inches and 3 inches, respectively, are melted and recast into a single disk having a diameter of 6 inches. Find its thickness.
Answer by mananth(12270)

1/4 inch - 5 " diameter = 2.5 inch radius.
volume = pi*2.5*2.5*1/4
V= 4.91 cu.in
1/6 inch -3 "diamter = 1.5" radius
V= pi*1.5*1.5*1/6
= 1.18 cu.in
Total = 4.91+1.18=6.09
...
Volume = pi*3*3*h
6.09/(pi*3*3)=h
h= 0.215inch or 1/5"
...
m.ananth@hotmail.ca